# Comprehend, Divide, and Conquer: Feature Subspace Exploration via Multi-Agent Hierarchical Reinforcement Learning

### v1

@@ Abstract.
ccs: Computing methodologies Feature selection
@@ 1. Introduction
Feature selection (FS) plays a critical role in classical machine learning models by mitigating the curse of dimensionality, reducing training times, and alleviating feature redundancy, thereby substantially enhancing predictive performance (li2017feature, ; wang2025towards, ; li2017recent, ). The evolution of FS methodologies has progressed from early heuristic and statistical screening techniques to the development of sophisticated approaches that incorporate filter-based (biesiada2008feature, ; yu2003feature, ; song2021fast, ), wrapper-based (stein2005decision, ; gheyas2010feature, ), and embedded strategies (dinh2020consistent, ; li2016deep, ; lemhadri2021lassonet, ) tailored to manage increasingly complex data environments. Concurrently, the rapid advancement of modern artificial intelligence has broadened the application domain of FS to encompass challenging tasks such as data pattern discovery (xiao2023beyond, ), biomarker identification (ying2024revolutionizing, ), and the construction of AI-driven data pipelines (AI4data) (liang2022advances, ). In the context of emerging demands in multi-modal data processing and real-time big data analytics, integrating efficient and intelligent feature selection is increasingly recognized as a pivotal element in enhancing data quality and advancing data-centric paradigms(fu2024tabular, ).
@@
In summary, as depicted in Figure 1(a), classic FS methods perform feature selection through a single decision, which is efficient but analyses the nature of the features inadequately, resulting in lower performance. Present RL-based Methods as depicted in Figure 1(b). Those methods of subspace exploration come from wrapper feature selection ideas and have powerful feature selection capabilities. Still, they also have the inherent drawbacks of wrapper methods, i.e., they have $\mathcal{O}(N)$ time complexity and require a lot of time when dealing with large datasets. To mitigate the challenge of managing an overwhelming number of agents, group-based (fan2021autogfs, ) and interaction-wise (fan2020autofs, ) agent architectures have been proposed as promising solutions. Nevertheless, these approaches rely solely on superficial mathematical characteristics (e.g., standard variance of features) to organize features, resulting in inaccuracies due to neglecting semantic and contextual feature relationships (li2024exploring, ; moraffah2024causal, ).
Those observations motivate us to develop novel hierarchical reinforcement learning-based architectures for feature sub-space exploration. The core concept of this research is that when a feature (e.g., age) is irrelevant to a task, its semantically equivalent feature (e.g., birthdate) may also be irrelevant. Additionally, two features exhibiting similar numerical patterns (e.g., highly related biological signals) could be redundant, necessitating fine-grained differentiation. Addressing both challenges demands an in-depth comprehension of feature traits and an innovative design of RL agent architecture.
@@ 2. Important Definitions
Feature Selection. Given the dataset $D=\{X\in\mathcal{R}^{m\times n},y\in\mathcal{R}^{m\times 1},F\}$, where $X$ and $y$ are the features and labels, respectively. $m$ and $n$ is the number of samples and features. We use a finite set $F=\{f_{1},f_{2},...,f_{n}\}$ to indicate the feature column included in dataset $D$, where $f_{i}$ is the $i-th$ column of $X$. The goal of feature selection is to find the optimal feature subset $F^{*}\subset F$ to enhance model performance while maintaining computational efficiency.
@@ 3.3. Divide-and-Conquer Agent Architecture
2:  $C_{i}\leftarrow\{f_{i}\}$
3:  $\mathcal{C}\leftarrow\ \{C_{1},...,C_{n}\}$
4:  $\mathcal{C}_{\text{agent}}\leftarrow\ \{\}$
@@
7:     $d_{ij}=\frac{|C_{i}||C_{j}|}{|C_{i}|+|C_{j}|}\|\overline{h_{ci}}-\overline{h_{cj}}\|^{2}$, where ||$\cdot$|| ⋅ | represents the number of items and $\overline{h_{ci}},\overline{h_{cj}}$ is the mean vector of each feature’s state within $C_{i}$ and $C_{j}$.
@@
$\diamondsuit$ State The state for each agent is a vectorized representation derived from the selected feature subset. We adopt the hybrid feature state as the state representation for each agent. We first acquire the parameter $\theta^{*}_{i}=\{z^{i}_{1},\cdots,z^{i}_{k},\mu^{i}_{1},\cdots,\mu^{i}_{k},\sigma^{i}_{1},\cdots,\sigma^{i}_{k}\}$ from GMM for the feature $f_{i}$. Due to the $z$ adjustment of the weight of each Gaussian component, we weighted-sum each $\mu$ and $\sigma$ toward corresponding $z$ to form a state vector $s_{i}$, formally:
@@
where $F$ is the entire feature set, $F_{t}$ is the selected feature subset in step-$t$, $\lambda$ is a hyperparameter, and ||$\cdot$|| ⋅ | denoted the size of given set. As $\lambda$ increases, the penalty for keeping too many features (large $|\mathcal{F}_{t}|$) becomes more severe, thus encouraging more substantial gene reduction. Conversely, a lower value of $\lambda$ relaxes the penalty against the size of $|\mathcal{F}_{t}|$, suitable when minimal reduction is sufficient. This metric ensures that the selection process strategically reduces the number of features.
@@ 3.4. Exploration and Opitmization
Analysis of the Advancement of HRL. The hierarchical agent structure activates fewer agents in each step, leading to superior performance in the iteration. The single-agent approach uses one agent to make $N$ decisions on $N$ features. Meanwhile, the multi-agent approach uses $N$ agents to make $N$ decisions on $N$ features, and the time complexity of both approaches is $\mathcal{O}(N)$. In contrast, we apply a hierarchical structure, which reduces the number of activated agents making decisions, totally using $12N-$ 1 agents to make $log(N)$ decisions on $N$ features, with time complexity $\mathcal{O}(log(N))$. In the following section, we will provide detailed proof of the time complexity advantage.
@@ 3.5. The Proof of Time Complexity of Comprehend-Divide-and-Conquer Solution
(21) $\begin{split}E(N)&=1+2pE(N/2)\\ &=1+2p[1+2pE(N/4)]\\ &=\sum_{k=0}^{h-1}(2p)^{k}+(2p)^{h}E(1)\\ &=\frac{(2p)^{h}-1}{2p-1}+(2p)^{h}\cdot 1\\ &=\frac{(2p)^{h+1}-1}{2p-1}\\ \end{split}$
@@ 4.3. Baseline Algorithms
We conducted a comparison with 8 methods, encompassing both classical and recent approaches. The baseline methods include: KBest (yang1997comparative, ) selects the K-best features based on statistical measure. MCDM (mcdm, ) uses predefined rankers to assign scores to features and select them accordingly. LASSONet (lemhadri2021lassonet, ) utilizes a residual component that integrates feature selection with the model training process, enhancing its generalizability. CompFS (yasuda2023sequential, ) provides an efficient, single-pass implementation of greedy forward selection, utilizing attention weights at each step to approximate feature importance. SAFS (imrie2022composite, ) ensembles feature selection models within a neural network to identify feature groups and measures the similarity between these groups to select the optimal feature subset. GAINS (xiao2023beyond, ) embeds features into a vector space and identifies the optimal subset using a gradient ascent search algorithm. SARLFS (sarlfs, ) constructs an agent to determine the subset of features. MARLFS (marlfs, ) is a multi-agent system in which the number of agents matches the number of features.
@@ 4.4. Hyperparameter Settings and Reproducibility
We performed 200 epochs for HRLFS to explore the feature space and an additional 200 epochs to optimize the agents, with the memory size set to 400. Follow the hyperparameter and ablation experiment results, The policy model of all agents are Actor-Critic, where the actor and critic models are both implemented as two-layer neural networks, with (64, 8) as hidden size. We employed Adam to optimize actor and critic models with a minibatch size of 32. The learning rates of the actor and critic models were 0.001 and 0.01 respectively. The hyperparameter $\gamma$ in Equation 8 as 0.9. We adopted Random Forest as the downstream machine learning model to evaluate the performance of the selected feature subset in each step. We set all features’ Gaussian component number $k$ as the maximum number of their BIC search result. For feature metadata embedding, we apply text-embedding-3-large (openai_embedding, ) based on GPT-4 (openai_gpt4, ). We set the hyperparameters $\alpha$ in Equation 6 and $\lambda$ in Equation 5 as 0.4 and 0.6 based on the results from the hyperparameter study in section 5.8. The hyperparameter settings for the baselines follow their original articles.
@@ 5. Experimental Results
Dataset Task #Samp. #Feat. KBest LASSONet MCDM SARLFS MARLFS SAFS GAINS CompFS Ours
@@
SVMGuide3 C 1243 21 76.55 77.61 75.68 75.41 75.92 75.53 77.53 72.83 $\mathbf{78.90^{\pm 0.45}}$
German_Credit C 1001 24 65.75 58.85 70.49 61.31 66.89 66.89 69.40 68.27 $\mathbf{73.07^{\pm 1.13}}$
Credit_Default C 30000 25 80.40 79.86 74.87 80.28 80.11 77.49 80.05 78.15 $\mathbf{80.56^{\pm 0.09}}$
SpamBase C 4601 57 90.36 89.16 89.10 91.28 90.04 90.93 91.71 89.70 $\mathbf{92.60^{\pm 0.31}}$
Megawatt1 C 253 38 84.02 89.59 89.59 88.02 87.64 86.46 89.42 91.83 $\mathbf{92.70^{\pm 0.73}}$
Ionosphere C 351 34 91.48 90.17 88.85 92.82 90.18 90.18 92.24 91.08 $\mathbf{94.27^{\pm 1.04}}$
Mice-Protein MC 1080 77 82.86 82.71 81.95 79.16 80.56 77.81 81.04 78.96 $\mathbf{83.79^{\pm 0.91}}$
Coil-20 MC 1440 400 93.92 88.19 96.53 96.88 96.19 94.41 97.22 95.16 $\mathbf{97.56^{\pm 0.06}}$
MNIST MC 10000 784 89.67 86.40 91.25 89.10 90.75 86.65 91.40 87.65 $\mathbf{91.75^{\pm 0.24}}$
Otto MC 61878 93 72.31 71.65 73.24 71.95 73.74 72.13 72.40 71.97 $\mathbf{74.24^{\pm 0.07}}$
Jannis MC 83733 54 65.48 65.78 64.20 66.05 66.70 64.42 65.15 66.61 $\mathbf{67.91^{\pm 0.22}}$
Cao MC 4186 13488 82.81 83.65 82.37 85.57 83.71 84.67 82.47 85.57 $\mathbf{89.37^{\pm 0.13}}$
Han MC 2746 20670 70.18 72.57 71.33 75.68 74.69 75.79 74.03 73.20 $\mathbf{81.45^{\pm 0.21}}$
Openml_586 R 1000 25 56.35 56.03 55.82 51.93 55.02 53.15 59.36 55.49 $\mathbf{61.75^{\pm 0.71}}$
Openml_589 R 1000 25 53.34 52.99 54.12 36.36 52.68 45.51 59.33 48.46 $\mathbf{61.60^{\pm 1.01}}$
Openml_607 R 1000 50 53.63 53.70 55.37 55.61 54.82 51.82 60.44 53.18 $\mathbf{61.70^{\pm 0.69}}$
Openml_616 R 500 50 31.16 21.69 28.27 27.50 31.36 46.16 49.54 49.16 $\mathbf{53.90^{\pm 0.86}}$
Openml_618 R 1000 50 49.21 48.97 48.69 48.03 49.31 45.16 55.68 48.55 $\mathbf{56.40^{\pm 0.62}}$
Openml_620 R 1000 25 53.00 54.63 55.00 35.95 54.57 50.16 61.51 55.49 $\mathbf{62.78^{\pm 0.82}}$
Openml_637 R 500 50 23.09 25.42 23.75 23.06 25.96 34.18 39.46 36.17 $\mathbf{40.54^{\pm 0.37}}$
@@ 5.1. Main Comparison
This experiment aims to answer the question: Is HRLFS capable of effectively refining a superior feature subspace across all domains? We compared our method with classical and recent approaches across various datasets. Table 1 shows the overall results of HRLFS. We observed that HRLFS outperforms the classical, deep learning-based, and RL-based methods on all tasks. The underlying driver is that, compared to the classical methods (KBest, LASSONet, and MCDM), HRLFS discovers fine-grained differentiation between features through LLMs to construct a more accurate feature subspace beyond numerical statistics, resulting in better performance. Compared to deep learning-based methods (SAFS, CompFS, GAINS), HRLFS manages clustered features using a hierarchical architecture that considers the correlation between features from coarse to fine grain, enables better exploration of feature subspace. Compared to RL-based methods (SARLFS, MARLFS), agent collaboration is viewed through the hierarchical reinforcement learning architecture in HRLFS, leading to an improvement in performance. Overall, this experiment illustrates that HRLFS is prominent and robust across diverse datasets, emphasizing various tasks, sizes, and fields, underlining its universal applicability for feature subspace exploration.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Feature selection
@@ 1. Introduction
Feature selection (FS) plays a critical role in classical machine learning models by mitigating the curse of dimensionality, reducing training times, and alleviating feature redundancy, thereby substantially enhancing predictive performance (li2017feature, ; wang2025towards, ; li2017recent, ). The evolution of FS methodologies has progressed from early heuristic and statistical screening techniques to the development of sophisticated approaches that incorporate filter-based (biesiada2008feature, ; yu2003feature, ; song2021fast, ), wrapper-based (stein2005decision, ; gheyas2010feature, ), and embedded strategies (dinh2020consistent, ; li2016deep, ; lemhadri2021lassonet, ) tailored to manage increasingly complex data environments. Concurrently, the rapid advancement of modern artificial intelligence (qin2025scihorizon, ; liugut, ) has broadened the application domain of FS to encompass challenging tasks such as data pattern discovery (xiao2023beyond, ), biomarker identification (11164312, ; ying2024revolutionizing, ), and the construction of AI-driven data pipelines (AI4data) (liang2022advances, ; cai2025knowledge, ; huang2025collaborative, ; xiao2025m, ). In the context of emerging demands in multi-modal data processing and real-time big data analytics, integrating efficient and intelligent feature selection is increasingly recognized as a pivotal element in enhancing data quality and advancing data-centric paradigms (fu2024tabular, ).
@@
In summary, as depicted in Figure 1(a), classic FS methods perform feature selection through a single decision, which is efficient but analyzes the nature of the features inadequately, resulting in lower performance. Present RL-based Methods as depicted in Figure 1(b). Those methods of subspace exploration come from wrapper feature selection ideas and have powerful feature selection capabilities. Still, they also have the inherent drawbacks of wrapper methods, i.e., they have $\mathcal{O}(N)$ time complexity and require a lot of time when dealing with large datasets. To mitigate the challenge of managing an overwhelming number of agents, group-based (fan2021autogfs, ) and interaction-wise (fan2020autofs, ) agent architectures have been proposed as promising solutions. Nevertheless, these approaches rely solely on superficial mathematical characteristics (e.g., standard variance of features) to organize features, resulting in inaccuracies due to neglecting semantic and contextual feature relationships (li2024exploring, ; moraffah2024causal, ).
[…]
@@ 2. Important Definitions
Feature Selection. Given the dataset $D=\{X\in\mathcal{R}^{m\times n},y\in\mathcal{R}^{m\times 1},F\}$, where $X$ and $y$ are the features and labels, respectively. $m$ and $n$ is the number of samples and features. We use a finite set $F=\{f_{1},f_{2},...,f_{n}\}$ to indicate the feature column included in dataset $D$, where $f_{i}$ is the ithi-th column of $X$. The goal of feature selection is to find the optimal feature subset $F^{*}\subset F$ to enhance model performance while maintaining computational efficiency.
@@ 3.3. Divide-and-Conquer Agent Architecture
[…]
@@
7:  $d_{ij}=\frac{|C_{i}||C_{j}|}{|C_{i}|+|C_{j}|}\|\overline{h_{ci}}-\overline{h_{cj}}\|^{2}$, where $|\cdot|$ represents the number of items and $\overline{h_{ci}},\overline{h_{cj}}$ is the mean vector of each feature’s state within $C_{i}$ and $C_{j}$.
@@
$\diamondsuit$ State The state for each agent is a vectorized representation derived from the selected feature subset. We adopt the hybrid feature state as the state representation for each agent. We first acquire the parameter $\theta^{*}_{i}=\{z^{i}_{1},\cdots,z^{i}_{k},\mu^{i}_{1},\cdots,\mu^{i}_{k},\sigma^{i}_{1},\\ \cdots,\sigma^{i}_{k}\}$ from GMM for the feature $f_{i}$. Due to the $z$ adjustment of the weight of each Gaussian component, we weighted-sum each $\mu$ and $\sigma$ toward the corresponding $z$ to form a state vector $s_{i}$, formally:
@@
where $F$ is the entire feature set, $F_{t}$ is the selected feature subset in step-$t$, $\lambda$ is a hyperparameter, and $|\cdot|$ denoted the size of given set. As $\lambda$ increases, the penalty for keeping too many features (large $|\mathcal{F}_{t}|$) becomes more severe, thus encouraging more substantial gene reduction. Conversely, a lower value of $\lambda$ relaxes the penalty against the size of $|\mathcal{F}_{t}|$, suitable when minimal reduction is sufficient. This metric ensures that the selection process strategically reduces the number of features.
@@ 3.4. Exploration and Opitmization
Analysis of the Advancement of HRL. The hierarchical agent structure activates fewer agents in each step, leading to superior performance in the iteration. The single-agent approach uses one agent to make $N$ decisions on $N$ features. Meanwhile, the multi-agent approach uses $N$ agents to make $N$ decisions on $N$ features, and the time complexity of both approaches is $\mathcal{O}(N)$. In contrast, we apply a hierarchical structure, which reduces the number of activated agents making decisions, totally using 2N12N-1 agents to make $log(N)$ decisions on $N$ features, with time complexity $\mathcal{O}(log(N))$. In the following section, we will provide detailed proof of the time complexity advantage.
@@ 3.5. The Proof of Time Complexity of Comprehend-Divide-and-Conquer Solution
While this assumption is theoretically convenient, in practice our hierarchy closely approximates such a structure. We provide an empirical validation of this assumption in Section 5.9, showing that the balance factors and tree heights observed across 21 datasets are strongly aligned with those of a perfect binary tree.
@@
[…]
@@
In Section 5.2, we conducted experiments and the results show that the active agents decreased by 70.61% to 82.30%, and time consumption reduced by 35.49% to 55.26%. These observations are aligned with the conclusion.
@@ 4.3. Baseline Algorithms
We conduct extensive comparisons against 8 methods, encompassing both classical and recent approaches, covering three categories of feature selection: filter, embedded, and RL-based (latest warpper methods). For filter methods, KBest (yang1997comparative, ) selects the K-best features based on statistical measures, and MCDM (mcdm, ) uses predefined rankers to assign scores to features and selects them accordingly. for embedded methods, LASSONet (lemhadri2021lassonet, ) utilizes a residual component that integrates feature selection with the model training process, enhancing its generalizability, and GAINS (xiao2023beyond, ) embeds features into a vector space and identifies the optimal subset using a gradient ascent search algorithm; for wrapper methods (including reinforcement learning-based ones), CompFS (yasuda2023sequential, ) provides an efficient, single-pass implementation of greedy forward selection, utilizing attention weights at each step to approximate feature importance, SAFS (imrie2022composite, ) ensembles feature selection models within a neural network to identify feature groups and measures the similarity between these groups to select the optimal feature subset. For RL-based methods, MARLFS (marlfs, ) is a multi-agent system in which the number of agents matches the number of features, and RLAS (RLAS, ) synergizes median-initialized filters, Q-learning-based module selection, and importance-aware random grouping.
@@ 4.4. Hyperparameter Settings and Reproducibility
We performed 200 epochs for HRLFS to explore the feature space and an additional 200 epochs to optimize the agents, with the memory size set to 400. Follow the hyperparameter and ablation experiment results, The policy model of all agents are Actor-Critic, where the actor and critic models are both implemented as two-layer neural networks, with (64, 8) as hidden size. We employed Adam to optimize actor and critic models with a minibatch size of 32. Following the existing research (wu2020finite, ), to ensure convergence, the critic’s learning rate should be around ten times greater than that for the actor. So the learning rates of the actor and critic models were 0.001 and 0.01, respectively. The hyperparameter $\gamma$ in Equation 8 as 0.9. We adopted Random Forest as the downstream machine learning model to evaluate the performance of the selected feature subset in each step. We set all features’ Gaussian component number $k$ as the maximum number of their BIC search result. For feature metadata embedding, we apply text-embedding-3-large (openai_embedding, ) based on GPT-4 (openai_gpt4, ). We set the hyperparameters $\alpha$ in Equation 6 and $\lambda$ in Equation 5 as 0.4 and 0.6 based on the results from the hyperparameter study in section 5.8. The hyperparameter settings for the baselines follow their original articles.
@@ 5. Experimental Results
Dataset Task #Samp. #Feat. KBest LASSONet MCDM RLAS MARLFS SAFS GAINS CompFS Ours
@@
SVMGuide3 C 1243 21 76.55 77.61 75.68 76.13 75.92 75.53 77.53 72.83 $\mathbf{78.90^{\pm 0.45}}$
German_Credit C 1001 24 65.75 58.85 70.49 64.91 66.89 66.89 69.40 68.27 $\mathbf{73.07^{\pm 1.13}}$
Credit_Default C 30000 25 80.40 79.86 74.87 80.41 80.11 77.49 80.05 78.15 $\mathbf{80.56^{\pm 0.09}}$
SpamBase C 4601 57 90.36 89.16 89.10 91.40 90.04 90.93 91.71 89.70 $\mathbf{92.60^{\pm 0.31}}$
Megawatt1 C 253 38 84.02 89.59 89.59 83.22 87.64 86.46 89.42 91.83 $\mathbf{92.70^{\pm 0.73}}$
Ionosphere C 351 34 91.48 90.17 88.85 94.11 90.18 90.18 92.24 91.08 $\mathbf{94.27^{\pm 1.04}}$
Mice-Protein MC 1080 77 82.86 82.71 81.95 77.78 80.56 77.81 81.04 78.96 $\mathbf{83.79^{\pm 0.91}}$
Coil-20 MC 1440 400 93.92 88.19 96.53 90.97 96.19 94.41 97.22 95.16 $\mathbf{97.56^{\pm 0.06}}$
MNIST MC 10000 784 89.67 86.40 91.25 88.30 90.75 86.65 91.40 87.65 $\mathbf{91.75^{\pm 0.24}}$
Otto MC 61878 93 72.31 71.65 73.24 72.86 73.74 72.13 72.40 71.97 $\mathbf{74.24^{\pm 0.07}}$
Jannis MC 83733 54 65.48 65.78 64.20 57.88 66.70 64.42 65.15 66.61 $\mathbf{67.91^{\pm 0.22}}$
Cao MC 4186 13488 82.81 83.65 82.37 90.57 83.71 84.67 82.47 85.57 $\mathbf{89.37^{\pm 0.13}}$
Han MC 2746 20670 70.18 72.57 71.33 74.73 74.69 75.79 74.03 73.20 $\mathbf{81.45^{\pm 0.21}}$
Openml_586 R 1000 25 56.35 56.03 55.82 51.04 55.02 53.15 59.36 55.49 $\mathbf{61.75^{\pm 0.71}}$
Openml_589 R 1000 25 53.34 52.99 54.12 49.36 52.68 45.51 59.33 48.46 $\mathbf{61.60^{\pm 1.01}}$
Openml_607 R 1000 50 53.63 53.70 55.37 57.42 54.82 51.82 60.44 53.18 $\mathbf{61.70^{\pm 0.69}}$
Openml_616 R 500 50 31.16 21.69 28.27 50.38 31.36 46.16 49.54 49.16 $\mathbf{53.90^{\pm 0.86}}$
Openml_618 R 1000 50 49.21 48.97 48.69 58.57 49.31 45.16 55.68 48.55 $\mathbf{56.40^{\pm 0.62}}$
Openml_620 R 1000 25 53.00 54.63 55.00 41.63 54.57 50.16 61.51 55.49 $\mathbf{62.78^{\pm 0.82}}$
Openml_637 R 500 50 23.09 25.42 23.75 42.04 25.96 34.18 39.46 36.17 $\mathbf{40.54^{\pm 0.37}}$
@@ 5.1. Main Comparison
This experiment aims to answer the question: Is HRLFS capable of effectively refining a superior feature subspace across all domains? We compared our method with classical and recent approaches across various datasets. Table 1 shows the overall results of HRLFS. We observed that HRLFS outperforms the classical, deep learning-based, and RL-based methods on all tasks. The underlying driver is that, compared to the classical methods (KBest, LASSONet, and MCDM), HRLFS discovers fine-grained differentiation between features through LLMs to construct a more accurate feature subspace beyond numerical statistics, resulting in better performance. Compared to deep learning-based methods (SAFS, CompFS, GAINS), HRLFS manages clustered features using a hierarchical architecture that considers the correlation between features from coarse to fine grain, enables better exploration of feature subspace. Compared to RL-based methods (RLAS, MARLFS), agent collaboration is viewed through the hierarchical reinforcement learning architecture in HRLFS, leading to an improvement in performance. Overall, this experiment illustrates that HRLFS is prominent and robust across diverse datasets, emphasizing various tasks, sizes, and fields, underlining its universal applicability for feature subspace exploration.